# Modeling Issues with Eye Tracking Data

### v1

@@ Abstract
I describe and compare procedures for binary eye-tracking (ET) data. These procedures are applied to both raw and compressed data. The basic GLMM model is a logistic mixed model combined with random effects for persons and items. Additional models address autocorrelation eye-tracking serial observations. In particular, two novel approaches are illustrated that address serial without the use of an observed lag-1 predictor: a first-order autoregressive model obtained with generalized estimating equations, and a recurrent two-state survival model. Altogether, the results of four different analyses point to unresolved issues in the analysis of eye-tracking data and new directions for analytic development.
@@ 2 Relevant Literature
The subject observations analyzed in the current study are eye-tracking fixations for visual targets, originally collected by [Ryskin2015]. The hierarchical structure of the data is characterized by trial-level data nested within subjects and items. For a visual presentation of the data structure, see Figure 3 of [Cho2018]. On each trial, the speaker and the listener participated together in a brief conversational exchange aimed at identifying a target object (the different objects constitute a set of items). Although they were seated in separate rooms, they communicated naturally through wireless microphones while their eye movements were recorded simultaneously with two remote eye trackers. The variable $t$ captures whether the speaker uniquely identified the target object or whether multiple objects in the listener’s visual scene could match the speaker’s descriptions. It reflects the amount of referential competition the listener must cognitively manage during real-time interpretation. Referential competition was hypothesized to make the task more difficult. The variable $Privileged$ reflects whether listeners could reduce referential competition based on information they infer from the speaker’s perspective; this variable was hypothesized to make the task easier.
@@ 3 Data
The data structure is organized around repeated measures, where eye-tracking observations are collected for subject–item–trial combinations. Each subject completed multiple trials, each associated with a visual display [Ryskin2015]. Each trial included two perspective-taking conditions ($t$ and $Privileged$) as previously described; these two conditions were assigned as the item level and thus varied across items. In the original data, observations consisted of $n=112$ observations at .01 second intervals. Data are recorded as the number of discrete instances in which a person is gazing at a target area (1) versus off-target (0). The total number of data points usually manifests as strings of 1s and 0s. The original data were presumably recorded at a frequency of 1000Hz and compressed to .01 second intervals. The resulting data file consisted of 995,232 records in long format for 152 subjects.
@@ 4 Modeling
The first model of the six models was run on the full uncompressed data set. I refer to this as the GLM model, which is given by:
@@
Here, the binary response $y_{ijt}$ for subject $i$ and item $j$ at time $t$ is modeled using a Bernoulli regression with fixed effects for t, Privileged, the interaction of Time with $t$ and $Privileged$, and random effects for subject and item. The LAG model was obtained by augmenting the GLM model with a lag-1 predictor $y_{ij,t-1}$ to account for serial correlation:
@@
The GLM model is altered more substantially by omitting random effects and incorporating a coefficient $\phi$ for residual lag-1 autocorrelation $\phi$ directly into the model specification rather than an observed lag variable. For this purpose, generalized estimating equations can be used. To show the unique relevance of this model to eye-tracking data, the method of GEE is expanded in some detail.
@@
Generalized Estimating Equations (GEE) [LiangZeger1986] provide a semiparametric approach for marginal regression with correlated data. For binary outcomes, let $p_{ijt}=\Pr(Y_{ijt}=1)$ denote the marginal mean for subject $i$ for item $j$ at time $t$. The mean model is written as
@@
The marginal variance of each Bernoulli observation is $\sigma_{ijt}^{2}=\sigma^{2}(Y_{ijt})=p_{ijt}(1-p_{ijt})$, and the variance-function matrix for cluster $ij$ is $A_{ij}=\mathrm{diag}(\sigma_{ij1}^{2},\ldots,\sigma_{ijT}^{2})$, where $T=112$. Serial or within-cluster correlation is introduced by specifying a working correlation matrix $R$ containing the single parameter $\phi$. For AR(1) structure, $R$ is a Toeplitz matrix with off-diagonal elements defined as $\phi^{|l-m|}$, where $l$ and $m$ are two time indices. The response vector $Y_{i}=(Y_{ij1},\ldots,Y_{ijT})^{\top}$ then has the variance structure:
$V_{ij}$ $=A_{ij}^{1/2}\,R\,A_{ij}^{1/2},$ (4)
$V_{ij}^{-1}$ $=A_{ij}^{-1/2}\,R^{-1}\,A_{ij}^{-1/2},$
Thus, $\phi$ enters the estimating equations through $R$.
@@
where $\hat{\phi}$ is the pooled autocorrelation, and the large-sample covariance of $\widehat{\beta}$ is estimated using the robust sandwich estimator is
@@
In Equations (5) and (7) it can be seen that $\phi$ affects both fixed-effect estimates and standard errors.
Several authors have noted that the GEE working covariance $V_{i}=A_{i}^{1/2}R_{\alpha}A_{i}^{1/2}$ plays a role analogous to the autocorrelation of latent residual processes in generalized linear mixed or state-space models for binary data. In particular, [ZegerLiangAlbert1988] demonstrated that serial correlation among binary observations can be represented either by specifying a latent Gaussian process with AR(1) correlation or by imposing the same AR(1) structure on the working correlation matrix in GEE. Related comparisons appear in [HeagertyZeger2000], who show that marginal models (GEE) and latent-process conditional models often share identical first- and second-moment structures, though they differ in higher-order behavior and interpretation.
Rather than controlling for autocorrelation with an observed lag-1 indicator, $\phi$ can me incorporated directly into the model’s variance structure. Residual structures are more likely to represent a stationary process, especially when $Time$ is included as a covariate. [Levin2018ChangeBlindness] previously applied a GEE model to analyze eye-tracking data; however, they used an unstructured correlation matrix as opposed to a first-order autoregressive structure. The AR1 variance structure achieves a higher degree of efficiency by estimating a single coefficient $\phi$, but this economy depends on the length of the time series.
@@ 6 Survival Modeling
RLE modeling of eye tracking responses is designed to model the 0-1 category frequencies, rather than the 0-1 data individually. This strategy was suggested previously in the discussion of Table 1. For this purpose, binomial modeling can be used in the framework of a GLMM. However, it is convenient to implement this analysis relative to runs within a subject-item cluster. In fact, the runs are treated as discrete instances of gaze that do not correspond to a linear process, either manifest or latent. For RLE data the following COX model was applied:
$K_{ijb}$ $\sim\text{Binomial}\!\left(n_{ijb},\,p_{ijb}\right),$ (8)
$\text{logit}\!\left(p_{ijb}\right)$ $=\mu+\beta_{1}\,\text{Privileged}_{ij}+\beta_{2}\,\text{Contrast}_{ij}$
$\quad+\beta_{3}\,\text{time}_{b}\!*\!\text{contrast}_{ij}+\beta_{4}\,\text{time}_{b}\!*\!\text{privileged}_{ij}$
$\quad+u_{i}+v_{j},$
where the random effects $u$ and $v$ share the same variance structure shown in Equation 2.
In this equation, the unit of analysis is run, where an individual time series has been compressed. For RLE compression, the block corresponds to $Y=0$ or $Y=1$. In this case, the RLE model is equivalent to modeling $y_{ijr}$ with weights $n_{ijr}$, corresponding to block length. After compression, each episode can be represented by an interval $Time_{\text{start}},Time_{\text{stop}}]$ and an associated event indicator that marks whether a transition to the opposite gaze state occurred at the end of the interval. This representation lends itself to an alternating two–state survival framework, where separate transition hazards are modeled for leaving the target $1\rightarrow 0$ and entering the target $0\rightarrow 1$ [kleinbaum2012survival, therneau2024survival]. Because RLE compression preserves the dwell structure, the resulting episode durations provide suitable survival intervals for Cox regression. The COX model is:
@@
The model incorporates a Cox proportional hazards framework, the process being modeled is not survival in the conventional sense. The data consist of a binary time series in which each subject–item trajectory alternates repeatedly between two states, with each interval representing the duration spent in a given state before switching to the other. In this context, the COX model is used as a flexible estimator of state-specific switching intensities rather than time to a terminal event. By stratifying on transition direction, we allow separate baseline hazards for $0\rightarrow 1$ and $1\rightarrow 0$ transitions, so that covariate effects are interpreted as changes in the instantaneous rate of leaving the current state.
@@ 7 Results
Analyses of the GLM, LAG, RLE, and COX methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function geeglm from the package geepack [Halekoh2006_geepack] was used for analysis of the AR1 model, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival].
Because of these data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was structurally missing for the last run. To account for nonindependence across, and subjects-by-item cluster indicator was used as the cluster variable in both the AR1 and COX models.
Table 2: Fixed-effect estimates for four methods.
Term GLM AR1 LAG COX
Intercept -0.437 -0.393 -2.946
Privileged 0.077 0.052 0.083 0.053
Contrast -0.079 -0.057 -0.013 -0.046
Time 0.854 0.810 0.498
Priv*Time -0.012 0.000 0.012
Contrast*Time -0.015 -0.005 0.028
@@ 7.1 Fixed Effects
The estimates of fixed effect for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and $t$. It can be seen that the GLM and methods provide estimates dissimilar to the AR1 and COX methods, which are relatively similar to each other. The LAG model incorporating the lag-1 $y$ predictor stands out. In particular, the coefficients for $Time$ and $t$ are smaller, and the coefficient for $Privileged$ is higher. Note the variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. Also, the COX method did not produce intercepts or time interactions.
The fact that the AR1 and LAG models had large disparities may be due to the difference between modeling observed and residual autocorrelations. Both the AR1 and COX approaches employ theoretical modeling approaches to autocorrelation. the AR1 model incorporated $\phi$ for serial correlation, and the survival model incorporates features for phase transition. The GLM and AR1 models also indicate negligible interactions over time for the experimental variables. Finally, the GLM, AR1, and LAG models indicate that over time, the likelihood of being on-target increases.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide dissimilar SEs. The GLM provides the lowest, and the AR1 and COX models provide the highest. The LAG model produces SEs about double of those for the GLM model. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator.
Table 3: Standard errors of fixed-effect estimates for four methods.
Term GLM AR1 LAG COX
Intercept 0.035 0.014 0.039
Privileged 0.007 0.034 0.014 0.030
Contrast 0.004 0.021 0.009 0.018
Time 0.003 0.014 0.006
Priv*Time 0.007 0.032 0.015
Contrast*Time 0.005 0.020 0.009
@@
7.3 Random Effects
Random effects are given in Table 4 for all models except AR and the COX model. For the COX model, it is possible to include random effects, but the cluster option was used in lieu of adding random effects. The estimated variances obtained for the GLM and LAG methods were similar. For the AR1 model, the estimated autocorrelation is extremely high $\phi=.95$. However, this is not unexpected given an average run length of over 40. This value does raise the concern that this degree of autocorrelation may destabilize other coefficients, yet the moderate degree of similarity of the experimental effects and SEs for the AR1 and COX models suggests stable estimation.
@@ 7.3 Random Effects
Model Subject variance Item variance AR(1) $\alpha$ Scale
@@
AR1 0.948 0.993
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, AR1 and COX provide more conservative alternatives for estimating fixed effects. In the case of the AR1, this appears to result from modeling autocorrelation and avoiding the use of random effects. The results of the AR1 and COX models raise important and unresolved questions. First, the finding that these approaches provide very similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions. Of course, the similarity of the AR1 and COX models does not imply they are correct. Rather it raises the question of what should be modeled. Transitions are one choice, the probability of being in state 1 (on-target) at each time is another. The AR1 and COX models appears incorporate the first, and GLM and LAG focus on the second.
@@
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. Both the coefficient $\phi$ and the group-level random intercepts can account for the same temporal structure in the data. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017]. Under such conditions, this may lead to estimation instability [jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint use of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000].
@@
Data pre-processing is another important issue. Eye-tracking systems commonly compress data to avoid creating mammoth output files. Yet even with preliminary machine compression, files may be large. The RLE encoding for additional compression is convenient for survival analysis. Pre-processing also includes data cleaning which involves tricky decisions of which experimental records to retain for analysis. This is especially the case in eye-tracking studies because momentary distractions can result in several consecutive seconds of invalid responses.
@@
The results of this study collectively support the existence of experimental effects for the variables $Privileged$ and $t$. However, for the AR1 and COX models, the effect for $Privileged$ is marginally significant. The interaction terms for GLM and AR1 indicate the effects of $Privileged$ and $t$ are uniform over time. Different estimates of main effects in the current study varied in statistical strength which demonstrates that modeling assumptions can make a substantive different in both effects size and statistical significance. THe current study is based on a single data set, yet it does raise copnceptual issues in the statistical modeling of eye-tracking data. This provides a useful direction for future research studies.

### v3

@@ Abstract
I describe and compare procedures for binary eye-tracking (ET) data. These procedures are applied to both raw and compressed data. The basic GLMM model is a logistic mixed model combined with random effects for persons and items. Additional models address autocorrelation eye-tracking serial observations. In particular, two novel approaches are illustrated that address serial without the use of an observed lag-1 predictor: a first-order autoregressive model obtained with generalized estimating equations, and a recurrent two-state survival model. Altogether, the results of five different analyses point to unresolved issues in the analysis of eye-tracking data and new directions for analytic development.
@@ 2 Relevant Literature
The subject observations analyzed in the current study are eye-tracking fixations for visual targets, originally collected by [Ryskin2015]. The hierarchical structure of the data is characterized by trial-level data nested within subjects and items. For a visual presentation of the data structure, see Figure 3 of [Cho2018]. On each trial, the speaker and the listener participated together in a brief conversational exchange aimed at identifying a target object (the different objects constitute a set of items). Although they were seated in separate rooms, they communicated naturally through wireless microphones while their eye movements were recorded simultaneously with two remote eye trackers. The variable $Contrast$ captures whether the speaker uniquely identified the target object or whether multiple objects in the listener’s visual scene could match the speaker’s descriptions. It reflects the amount of referential competition the listener must cognitively manage during real-time interpretation. Referential competition was hypothesized to make the task more difficult. The variable $Privileged$ reflects whether listeners could reduce referential competition based on information they infer from the speaker’s perspective; this variable was hypothesized to make the task easier.
@@ 3 Data
The data structure is organized around repeated measures, where eye-tracking observations are collected for subject–item–trial combinations. Each subject completed multiple trials, each associated with a visual display [Ryskin2015]. Each trial included two perspective-taking conditions ($Contrast$ and $Privileged$) as previously described; these two conditions were assigned as the item level and thus varied across items. In the original data, observations consisted of $n=112$ observations at .01 second intervals. Data are recorded as the number of discrete instances in which a person is gazing at a target area (1) versus off-target (0). The total number of data points usually manifests as strings of 1s and 0s. The original data were presumably recorded at a frequency of 1000Hz and compressed to .01 second intervals. The resulting data file consisted of 995,232 records in long format for 152 subjects.
@@ 4 Modeling
The first model was run on the full uncompressed data set. I refer to this as the GLM model, which is given by:
@@
Here, the binary response $y_{ijt}$ for subject $i$ and item $j$ at time $t$ is modeled using a Bernoulli regression with fixed effects for t, Privileged, the interaction of $Time$ with $Contrast$ and $Privileged$, and random effects for subject and item. The LAG model was obtained by augmenting the GLM model with a lag-1 predictor $y_{ij,t-1}$ to account for serial recursion:
@@
The GLM model is altered more substantially by omitting random effects and incorporating a coefficient $\phi_{A}$ as an alternative for the lag-1 autocorrelation. Error correlation is then incorporated directly into the variance structure. To show the application of this approach, generalized estimating equations can be used. To show the application of this structure, it is expanded in some detail.
@@
Generalized Estimating Equations (GEE) [LiangZeger1986] provide a semiparametric approach for marginal regression with correlated errors. For binary outcomes, let $p_{ijt}=\Pr(Y_{ijt}=1)$ denote the marginal mean for subject $i$ for item $j$ at time $t$. The mean model is written as
@@
The marginal variance of each Bernoulli observation is $\sigma_{ijt}^{2}=\sigma^{2}(Y_{ijt})=p_{ijt}(1-p_{ijt})$, and the variance-function matrix for cluster $ij$ is $A_{ij}=\mathrm{diag}(\sigma_{ij1}^{2},\ldots,\sigma_{ijT}^{2})$, where $T=112$. Serial or within-cluster correlation is introduced by specifying a working correlation matrix $R$ containing the single parameter $\phi_{A}$. For AR(1) structure, $R$ is a symmetric matrix with off-diagonal elements defined as $\phi_{A}^{|l-m|}$, where $l$ and $m$ are two time indices. The observed vector $Y_{i}=(Y_{ij1},\ldots,Y_{ijT})^{\top}$ then has the variance structure:
$V_{ij}$ $=\alpha A_{ij}^{1/2}\,R\,A_{ij}^{1/2},$ (4)
where $\alpha$ is the scale parameter. Thus, the residual correlation $\phi$ enters the estimating equations through $R$.
@@
where $\hat{\phi_{A}}$ is the residual correlation, and the large-sample covariance of $\widehat{\beta}$ is estimated using the robust sandwich estimator is
@@
In Equations (5) and (7) it can be seen that $\phi_{A}$ affects both fixed-effect estimates and standard errors.
Several authors have noted that the GEE working covariance $V_{i}=A_{i}^{1/2}R_{\alpha}A_{i}^{1/2}$ plays a role analogous to the autocorrelation of latent residual processes in GLM or state-space models for binary data. In particular, [ZegerLiangAlbert1988] demonstrated that serial correlation among binary observations can be represented either by specifying a latent Gaussian process with AR(1) correlation or by imposing the same AR(1) structure on the working correlation matrix in GEE. Related comparisons appear in [HeagertyZeger2000], who show that marginal models in GEE and latent-process conditional models often share identical first- and second-moment structures, though they differ in higher-order behavior and interpretation.
Rather than controlling for autocorrelation with an observed lag-1 indicator, $\phi_{A}$ can me incorporated directly into the model’s variance structure. Residual structures are more likely to represent a stationary process, especially when $Time$ is included as a covariate. [Levin2018ChangeBlindness] previously applied a GEE model to analyze eye-tracking data; however, they used an unstructured correlation matrix as opposed to a first-order error structure. The AR1 variance structure achieves a higher degree of efficiency by estimating a single coefficient $\phi_{A}$, but this economy depends on the length of the time series.
A related model to AR1 is a first-order moving average model MA1. When binary data is composed of runs (streaks of 0s and 1s), it suggests that once a saccade occurs (e.g., a state switch), its influence is immediate but might not persist indefinitely across the entire time series. This seems more natural to a eye-tracking time series than an autogressive process. This model can also be fit with GEE in the forms of an $R$ matrix having a tri-diagonal structure, also known as a banded Toeplitz matrix of width 1. For this correlation matrix, only the off-diagonals immediately above and below the main diagonal have a nonzero value of $\phi_{M}$. The estimating equations are identical to those of the AR1 model.
@@ 6 Survival Modeling
RLE modeling of eye tracking responses is designed to model the 0-1 category frequencies, rather than the 0-1 data individually. This strategy was suggested previously in the discussion of Table 1. For this purpose, binomial modeling can be used in the framework of a GLMM. However, it is convenient to implement this analysis relative to runs within a subject-item cluster. In fact, the runs are treated as discrete instances of gaze that do not correspond to a linear process, either manifest or latent.
For RLE compression, the blocks correspond to $Y=0$ or to $Y=1$. After compression, each episode can be represented by an interval $Time_{\text{start}},Time_{\text{stop}}]$ and an associated event indicator that marks whether a transition to the opposite gaze state occurred at the end of the interval. This representation lends itself to an alternating two–state survival framework, where separate transition hazards are modeled for leaving the target $1\rightarrow 0$ and entering the target $0\rightarrow 1$ [kleinbaum2012survival, therneau2024survival]. Because RLE compression preserves the dwell structure, the resulting episode durations provide suitable survival intervals for Cox regression. The COX model is:
@@
The model incorporates a Cox proportional hazards framework, the process being modeled is not survival in the conventional sense. The data consist of a binary time series in which each subject–item trajectory alternates repeatedly between two states, with each interval representing the duration spent in a given state before switching to the other. In this context, the COX model is used as a flexible estimator of state-specific switching intensities rather than time to a terminal event. By stratifying on transition direction, separate baseline hazards arfe used for $0\rightarrow 1$ and $1\rightarrow 0$ transitions, so that covariate effects are interpreted as changes in the instantaneous rate of leaving the current state.
@@ 7 Results
Analyses of the GLM and LAG methods were carried out with the R function glmer from the package lme4 [Bates2015_lme4]. The R function glmgee from the R package glmgee [Vanegas2023] was used for analysis of the AR1/MA1 models, and the R function coxph from the package survival was used for the two-state COX model [therneau2024survival]. The AR1 and MA1 models produced positive definite $R$ matrices; however, their determinants were near zero. To address this issue, user-defined $R$ matrices were constructed with $\phi=.95$ and a ridge of 1e-5 was applied to $R$. In both cases, estimates were nearly identical to those produced in runs where $\phi$ was freely estimated.
Because of the data exclusion procedures used in the present study, analytic results cannot be directly compared to prior studies. In this section, only the models are evaluated. The data for the models was modified in two ways. First, models using a lag-1 predictor require dropping the first observation in each individual time series. Second, for the survival model, the last run was deleted. Because the model focuses on transitions, the transition indicator was missing by fiat for the last run. To account for non-independence across subject and items, and a subjects-by-item indicator was used as a cluster variable in the AR1, MA1, and COX models.
Table 2: Fixed-effect estimates for five methods.
Term GLM AR1 MA1 LAG COX
Intercept -0.437 -0.393 -0.400 -2.946
Ylag-1 5.784
Privileged 0.077 0.052 0.070 0.083 0.053
Contrast -0.079 -0.057 -0.068 -0.013 -0.046
Time 0.854 0.813 0.831 0.498
Priv*Time -0.012 0.001 0.006 0.012
Contrast*Time -0.015 -0.005 -0.012 0.028
@@ 7.1 Fixed Effects
The estimates of fixed effects for these models are given in Table 2. For the COX analysis, the total effect is given which consists of the aggregate main effects of $Privileged$ and the transition interaction ($y=0\rightarrow y=1)$, and likewise for $Contrast$. It can be seen that the GLM and MA1 methods provide similar estimates of experimental effects, and the AR1 and COX methods also have similar estimates. The LAG model incorporating the lag-1 $y$ predictor stands out against the other four models. In particular, the LAG coefficients for $Time$ and $Contrast$ are smaller, and the coefficient for $Privileged$ is higher. A ContrastTimeContrast*Time interaction is also evident for the lag model, and the $Time$ coefficient is strikingly dissimilar to those of the other models. The GLM, AR1, and MA1 models indicate a strong positive relationship between time and the likelihood of being on-target. The variable $Time$ is implicitly contained in the survival analysis model through run start and stop times. For this reason, the COX method did not produce intercepts or time interactions.
It is interesting that the AR1 and MA1 models resulted in moderately different estimates of experimental effects. In this regard, the MA1 model is most similar to the GLM model. Recall that the AR1, MA1, and COX approaches employ variance modeling alternatives to error correlation, while no adjustment for autocorrelation is present in the GLM model. Both models also indicate negligible interactions over time for the experimental variables.
@@ 7.2 Standard Errors
Standard errors for the fixed effects are given in Table 3. The models provide notably dissimilar SEs. The GLM provides the lowest, and the AR1/MA1, COX models provide the highest for the experimental valriables. The LAG model produces SEs about double of those for the GLM model, but even so, LAG SEs are half the size those for AR1, MA1, and COX. As noted earlier, the COX model has a lower effective sample size because the last run in each individual time series is dropped due to the lack of a transition indicator. Still, the SEs for the AR1, MA1, and COX models are very similar. This may imply that these three models implicitly weight transition nieghborhoods higher than other areas of the time series.
Table 3: Standard errors of fixed-effect estimates for five methods.
Term GLM AR1 MA1 LAG COX
Intercept 0.035 0.014 0.016 0.039
Ylag-1 0.012
Privileged 0.007 0.034 0.037 0.014 0.030
Contrast 0.004 0.021 0.023 0.009 0.018
Time 0.003 0.014 0.015 0.006
Priv*Time 0.007 0.033 0.035 0.015
Contrast*Time 0.005 0.020 0.022 0.009
@@
7.3 Random Effects and $\phi$
Random effects are given in Table 4 for the GLM and LAG models. For the AR1, MA1, and COX models, the cluster option was used in lieu of adding random effects. The estimated variance components obtained for the GLM and LAG methods were similar. For the AR1 and MA1 model, the estimated residual correlations are extremely high. However, this is not unexpected given the median run length of 35 (out of 112). This level of $\phi$ for AR1 and MA1 raises the concern that other coefficients may be destabilized. This is also true for the LAG model because a large pseudo $R^{2}$ is implied by the high magnitude of the $Ylag1$ coefficient.
@@ 7.3 Random Effects and $\phi$
Model Subject variance Item variance $\phi$ $\alpha$
@@
AR1 0.95 (0.948) 0.993
MA1 0.95 (0.963) 1.001
  • Note. Ridge estimates are reported for AR1 and MA1 with $\phi=.95$. Estimated values of $\phi$ are given in parentheses.
@@ 8 Discussion
The results of this study may shed some light on the issue of autocorrelation in eye-tracking data. Relative to the GLM and LAG models, AR1/MA1 and COX provide more conservative alternatives for estimating fixed effects. The results of the AR1 and COX models raise important and unresolved questions. First, the finding that these two approaches provide similar estimates of experimental effects may not be a coincidence. Recent work has shown that binary time–series models with autoregressive structure can behave very much like discrete-time two–state transition or survival models. For example, [campajola2020ising] demonstrated that a binary autoregressive process is mathematically equivalent to a Markov two–state model. Similarly, [gorgi2016survival] demonstrated how autoregressive recurrences can be expressed in terms of survival or transition probabilities. These relationships may help to explain why fitting an autoregressive model to a binary sequence of blocks can produce coefficients that resemble those obtained from modeling $0\rightarrow 1$ transitions.
Of course, the similarity of the AR1 and COX models does not imply they are correct. Rather it raises the question of how eye-tracking time series can be modeled. The AR1 and MA1 models are impractical for longer time series because the determinant of $R$ approaches zero. While a upper boundary can be set manually, this is purely an ad hoc solution. The survival model, or similar models, may provide a way to avoid near singularities due to serial correlation. Moreover, these models are not truly time series models but rather error-correlation models. Based on results from the current study, it is not clear that time series models are require or even suitable for eye-tracking data.
@@
In hierarchical time-series models, it is well documented that autoregressive parameters and random effects are difficult to identify simultaneously. The coefficients $\phi$ of AR1/MA1 and $\beta_{0}$ of the LAG model can account for the same temporal structure as the subject and item random intercepts. When the available information is insufficient to distinguish their contributions, the model may distribute this dependence ambiguously between them, producing weak identification of $\phi$ and $\beta_{0}$, and unstable or inflated variance estimates for the random effects [diggle2002, galecki2013]. These difficulties arise particularly when individual time series are short [verbeke2000, shumway2017, jones1993]. For these reasons, methodological guidance commonly recommends avoiding the joint modeling of autocorrelation and random effects unless the study design provides strong and independent information to identify both components [pinheiro2000]. It is possible that a similar problem arises from modeling with a lag-1 dependent variable. For instance, in Table 2 the lag-1 coefficient seems to have redistributed $Time$ and $Contrast$ effects to the ContrastTimeContrast*Time interaction the subject variance component as shown in Table 4.
@@
Data pre-processing is another important issue. Eye-tracking systems commonly compress data to avoid creating mammoth output files. Yet even with preliminary machine compression, files may be large. The RLE encoding method for additional compression is convenient for survival analysis and greatly reduces execution time. Pre-processing also includes data cleaning which involves tricky decisions of which experimental records to retain for analysis. This is especially the case in eye-tracking studies because momentary distractions can result in several consecutive seconds of invalid responses.
@@
The results of this study collectively support the existence of experimental effects for the variable $Contrast$. For the AR1, MA1, and COX models, the effect for $Privileged$ is marginally significant. Also, the interaction terms for GLM and AR1/MA1 indicate the effects of $Privileged$ and $Contrast$ are uniform over time. Because estimates of the experimental effects in the current study varied in statistical strength, this demonstrates model assumptions can make substantive differences in both effect size and statistical significance.
The current study is based on a single data set, yet it does raise conceptual issues in the statistical modeling of eye-tracking data. Results are limited to a two-regime model for on-target and off-target, and a single data set. However, useful directions are suggested for future research on modeling issues in eye-tracking data. In particular, Markov transition models are suggested, but suitable software functions for the current data set would require capability for incomplete subject-by-item designs as well as clustering.